OpenStudy (anonymous):

Select the property illustrated in the statement. 5 + x = x + 5 Identity Property for Multiplication Inverse Property for Addition Commutative for Addition Distributive Property

OpenStudy (uri):

B

OpenStudy (anonymous):

it's B, right?

OpenStudy (uri):

Yes B.

OpenStudy (anonymous):

identity property of multiplication states a*1 = a inverse property of addition states A number added to its opposite integer is always equal to zero commutative property of addition states A + B = B + A distributive property for addition states (A+B)+C = A + (B+C) hope this helps :-)

OpenStudy (anonymous):

OHHHH it's C haha

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thanks :)